OpenStudy (anonymous):

solving linear systems using substitution method 2x+y=9 x+y=4

OpenStudy (anonymous):

2x+y=9 -2(x+y=4 so 2x+y=9 -2x-2y=-8 cross out the x's -1y=1 y=-1 then just use the first problem and put in the y to get the x and wuallah!

OpenStudy (anonymous):

i dont understand , can you draw it out ?

OpenStudy (anonymous):

i just did it lol

OpenStudy (anonymous):

after -2x+y=-8 i added +2 on both sides and got -2y over 2 + 10 over 2 and i came out with 10 over 2 equals 5